# Foldable SuperNets: Scalable Merging of Transformers with Different Initializations and Tasks

### v1

@@ 4 Conclusion
Limitations. FS-Merge requires a small unlabeled subset of the original training data, similarly to most previous merging methods. Additionally, FS-Merge, like distillation, is more computationally intensive than standard merging methods; however, these methods completely failed in more challenging settings. Notably, the main bottleneck in model merging is often data availability rather than computational resources. Moreover, FS-Merge has fewer learnable parameters than distillation, but they increase linearly with the number of models and hidden width; however, FS-Merge seq. solved the first issue. Lastly, one cannot naively merge two models of different depths using our method; we believe this could be solved in future work.
@@
Reproducibility. The paper fully discloses all the information needed to reproduce the experimental results. The method is detailed in Section 2; the main results are shown in Section 3; the pre-training details are explained in Appendix D; the full experimental details and hyperparameters are written in Appendix E; and the additional results can be found in Appendix F.1. In addition, code will be published upon acceptance.
Ethics. We can identify several aspects where FS-Merge can mitigate ethical concerns. The ability to merge models from different initializations and tasks offers an efficient alternative for using an ensemble of these models. This allows us to achieve greater resource efficiency and reduce the model’s carbon footprint. Moreover, FS-Merge can alleviate privacy concerns. For example, in cases where multiple users are training models on private datasets, FS-Merge enables us to combine those models into a single multi-task model, without accessing the full private datasets or any labels.
@@ A.1 Related work
A pair of models has mode connectivity when there exists a simple path between them in the loss landscape with low loss (garipov2018loss; draxler2018essentially; lubana2023mechanistic). frankle2020linear demonstrated that this path can be linear (LMC) when the models originate from the same initialization. tatro2020optimizing; ainsworth2023git; entezari2022the showed that it is feasible to establish LMC between MLPs and CNNs from different initializations using permutations. jordan2023repair enhanced these solutions by addressing their variance collapse. Wang2020Federated used these permutations to match and then merge models within a federated learning framework. imfeld2023transformer; verma2024merging used permutations to align two transformers with different initializations that were trained on the same task. mirzadeh2021linear leveraged LMC for multitask and continual learning. yunis2022on expanded LMC to more than two models, discovering a high-dimensional convex hull of low loss. zhou2023going introduced Layerwise Linear Feature Connectivity (LLFC), demanding that the features of the models be linearly connected. singh2020model; akash2022wasserstein addressed a similar challenge of neuron alignment using optimal transport (knight2008sinkhorn), and liu2022deep generalized it as a graph-matching task. navon2023equivariant solved the neuron alignments problem by training an equivariant deep weight space network. It is important to note that previous works which merged models from different initializations using permutations primarily focused on merging pairs of models trained on the same task, with many of them concentrating on MLPs and CNNs. In contrast, our method is capable of merging larger groups of models, with different initializations and tasks, and is specifically designed to handle transformers.
@@
Model merging technique (goddard2024arcee) has gained increasing interest in the past years, allowing the creation of stronger single-task and multi-task models. akhlaghi2018knowledge showed that averaging the weights of multiple simple neural networks maintains their performance. Model soups (wortsman2022model) proposed averaging multiple models trained on the same task from identical initializations to enhance task accuracy. In addition, Weight averaging has been employed for various purposes, including improving optimization through checkpoint averaging (izmailov2018averaging), federated learning (mcmahan2017communication), developing superior pre-trained models (choshen2022fusing; don2022cold), improving Out-of-Distribution Generalization (rame2022diverse), achieving success in unseen tasks (huang2023lorahub), and as augmentations for weight space networks (shamsian2024improved). In many cases, spherical linear interpolation (SLERP) (shoemake1985animating) is used to average the models’ weights. matena2022merging utilized Fisher-Weighted Averaging to fuse multiple models from the same initializations but trained on diverse tasks, resulting in a multi-task model. RegMean (jin2023dataless) addressed a similar scenario and proposed a closed-form solution that solves a local linear regression problem for each linear layer in the model. ilharco2023editing defined task vector by subtracting the parameters of a fine-tuned model from those of the pre-trained model, and used it to fuse models fine-tuned from the same pre-trained model. yadav2024ties; ortiz2024task; yang2024adamerging; akiba2024evolutionary analyzed and proposed more merging methods based on task vectos. sung2023an fused pairs of models trained on different modalities. ZipIt (stoica2024zipit) merged models from various initializations and tasks, focusing on MLPs and CNNs, by averaging pairs of highly correlated neurons between and within the models (see Appendix A.2 for more details). In contrast, our work focuses on MLPs and transformers, and can use much more complicated merging schemes.
@@
In Knowledge Distillation (KD), a small student model is trained to mimic the outputs of a larger teacher model (ba2014deep; hinton2015distilling; stanton2021does; gou2021knowledge). sau2016deep implemented a noise-based methodology to simulate learning a single task from multiple teachers. tan2018multilingual; khanuja2021mergedistill used several teachers, each translating between a specific language pair, to train a singular multilingual student. clark2019bam; park2023multitask; chelaramani2021multi utilized multiple models and true labels to instruct a multi-task model. pham2023collaborative employed numerous quantized teachers trained on the same task to teach a quantized model, also leveraging true labels. wu2021one co-fine-tuned multiple teachers in downstream tasks with shared layers to instruct a student. jacob2023online simultaneously trained multiple single-task models with a single multi-task student to facilitate the student’s optimization. Similar to our distillation baseline, vongkulbhisal2019unifying utilized KD to unify knowledge from various models with different targets into a single model, relying solely on unlabeled data. wu2021one; zagoruyko2017paying; heo2019knowledge; heo2019comprehensive focused on using a single teacher’s outputs and inner features to train a single model. park2019feed; liu2020adaptive applied multiple teachers trained on the same task, their inner features, and the true labels to train a single-task student. Inspired these works, we tried to merge transformers using a loss function that takes into account the inner features. However, this approach proved ineffective for both our method and distillation. See Appendix H.3 for details.
@@ A.2 Comparison to ZipIt
Our work strongly relates to ZipIt (stoica2024zipit), which also merges models from various initializations and tasks, with a focus on MLPs and CNNs. We will briefly explain the algorithm and its limitations. ZipIt uses the original training data to extract features from the models to be merged, $A$ and $B$. Then, for each layer $l$, it employs the features to identify pairs of highly correlated neurons using a greedy algorithm. Following this, it builds $M_{l}\in\mathbb{R}^{d_{l}\times 2\cdot d_{l}}$, which is zero except for entries corresponding to a matched pair $(i,j)$ indexed by $p$, where $M_{l}[p,i]=M_{l}[p,j]=\frac{1}{2}$. The purpose of this matrix is to merge features by averaging pairs of highly correlated neurons. Then, $U_{l}=2M_{l}^{\top}$ is used to reposition the merged features back to their original locations. After building $M_{l}$ and $U_{l}$, ZipIt proposes folding them with the original weights (Eq. 5) to create the weights of the merged model.
Like other current model merging methods (ainsworth2023git; jin2023dataless; imfeld2023transformer), ZipIt is efficient and focuses only on simple fusion schemes. It limit itself to pairing similar neurons, addressing only the permutation symmetries of neural networks (hecht1990algebraic). Permutation symmetries mean that it is possible to swap any two neurons of a hidden layer in a neural network without altering its functionality. However, ZipIt falls short in handling the scale symmetries of neural networks (neyshabur2015path; badrinarayanan2015understanding; phuong2019functional), or in considering more complicated merging rules.
@@ B.1 Foldable SuperNet for Vision Transformers
Assuming there are two Vision Transformers (ViTs) (dosovitskiy2021an), trained on two distinct tasks, $A$ and $B$, we aim to define a Foldable SuperNet that combines the original weights of the ViTs with new learnable parameters $M,U$. This structure is designed so that all skip connections and layer normalizations (ba2016layer; xiong2020layer) operate on the merged dimension, and that there is a linear layer before or after every $M$ or $U$ matrix, allowing them to be folded after training. It is important to highlight that layer norms possess significantly fewer learnable parameters compared to other layers in ViTs. Therefore, we can initiate their parameters from a good starting point (e.g., the parameters of the first model to be merged) and proceed to optimize the parameters as usual, similar to strategies employed in previous merging works (jordan2023repair; stoica2024zipit).
@@ B.1.1 Pre-processing
The tokens from both models are then concatenated to form $(Z^{A}||Z^{B})\in\mathbb{R}^{T\times 2\cdot d}$, and the Foldable SuperNet merges them using a learned matrix ${\color[rgb]{1,0,0}M_{\text{in}}}\in\mathbb{R}^{2\cdot d\times d}$ to produce $Z^{*}\in\mathbb{R}^{T\times d}$.
$Z^{*}=(Z^{A}||Z^{B}){\color[rgb]{1,0,0}M_{\text{in}}}\,.$
Subsequently, and like the ViT, the Foldable SuperNet applies layer normalization, with parameters ${\color[rgb]{1,0,0}\gamma^{*}_{-1},\beta^{*}_{-1}}\in\mathbb{R}^{d}$, to generate the input for the transformer. These parameters will be optimized.
$\bar{X}_{0}^{*}=\text{LN}_{{\color[rgb]{1,0,0}\gamma^{*}_{-1},\beta^{*}_{-1}}}(Z^{*})$
@@ B.1.2 The attention block
The Foldable SuperNet of the attention block at layer $l$ (Figure  3) receives merged features $X^{*}_{l}\in\mathbb{R}^{T\times d}$ as inputs, and applies layer normalization. It’s parameters ${\color[rgb]{1,0,0}\gamma^{*}_{l},\beta^{*}_{l}}\in\mathbb{R}^{d}$ will be learned. Additionally, a learnable matrix ${\color[rgb]{1,0,0}U_{l}}\in\mathbb{R}^{d\times 2\cdot d}$ is introduced to reconstruct the original features from the merged ones.
$\tilde{X}_{l}^{A}||\tilde{X}_{l}^{B}=\text{LN}_{{\color[rgb]{1,0,0}\gamma^{*}_{l},\beta^{*}_{l}}}(X_{l}^{*}){\color[rgb]{1,0,0}U_{l}}\,.$
@@
$Q_{l}^{*}=(\tilde{Q}^{A}_{l}||\tilde{Q}^{B}_{l}){\color[rgb]{1,0,0}M_{l}^{Q}}=(Q_{1,l}^{*},...,Q_{H,l}^{*})\,.$
The weights $W_{l}^{Q,A},W_{l}^{Q,B}\in\mathbb{R}^{d\times d}$ generate queries from the embeddings. The matrix ${\color[rgb]{1,0,0}M_{l}^{Q}}\in\mathbb{R}^{2\cdot d\times d}$ merges these features, which are then divided into $H$ heads, where each head $Q_{1,l}^{*}$ has dimensions $\mathbb{R}^{T\times\frac{d}{H}}$. Similarly, the keys and values are created by $W_{l}^{K},W_{l}^{V}\in\mathbb{R}^{d\times d}$, and compressed using the matrices ${\color[rgb]{1,0,0}M_{l}^{K},M_{l}^{V}}\in\mathbb{R}^{2\cdot d\times d}$ respectively.
Following this, and similar to the ViT, the Foldable SuperNet executes multi-head attention with the merged queries, keys and values and concatenates the features from the heads. In our Foldable SuperNet, this step also includes adding ${\color[rgb]{1,0,0}U_{l}^{O}}\in\mathbb{R}^{d\times 2\cdot d}$ to reconstruct the original multi-head attention outputs. Then each ViT utilizes $W_{l}^{O}\in\mathbb{R}^{d\times d}$ to aggregate those outputs. We also use ${\color[rgb]{1,0,0}M_{l}^{O}}\in\mathbb{R}^{2\cdot d\times d}$ to compress it once more. This is followed by a skip connection.
$Y_{l}^{*}=X_{l}^{*}+\text{Concat}_{i}[...,\text{softmax}\left(\frac{Q_{i,l}^{*}{K_{i,l}^{*}}^{\top}}{\sqrt{d}}\right)V_{i,l}^{*},...]\ {\color[rgb]{1,0,0}U_{l}^{O}}\begin{pmatrix}W_{l}^{O,A}&0\\ 0&W_{l}^{O,B}\end{pmatrix}{\color[rgb]{1,0,0}M_{l}^{O}}\,.$

### v2

@@ 4 Discussion
Limitations. FS-Merge requires a small unlabeled subset of the original training data, similarly to most previous merging methods. Additionally, FS-Merge, like KD, is more computationally intensive than standard merging methods; however, these methods completely failed in more challenging settings. Moreover, FS-Merge has fewer learnable parameters than KD, but they increase linearly with the number of models and hidden width; however, FS-Merge seq. solved the first issue. Lastly, one cannot naively merge two models of different depths using our method; we believe this could be solved in future work.
@@
Impact Statement. FS-Merge addresses key challenges in AI development. The ability to merge models from different initializations and tasks offers an efficient alternative to using an ensemble of these models. This approach enables greater resource efficiency and reduces the model’s carbon footprint. Moreover, FS-Merge can alleviate privacy concerns. For instance, in cases where multiple users train models on private datasets, FS-Merge allows these models to be combined into a single multi-task model without accessing the full private datasets or any labels.
@@ A.1 Related work
A pair of models has mode connectivity when there exists a simple path between them in the loss landscape with low loss (Garipov et al., 2018; Draxler et al., 2018; Lubana et al., 2023). Frankle et al. (2020) demonstrated that this path can be linear (LMC) when the models originate from the same initialization. Tatro et al. (2020); Ainsworth et al. (2023); Entezari et al. (2022) showed that it is feasible to establish LMC between MLPs and CNNs from different initializations using permutations. Jordan et al. (2023) enhanced these solutions by addressing their variance collapse. Wang et al. (2020) used these permutations to match and then merge models within a federated learning framework. Imfeld et al. (2023); Verma & Elbayad (2024) used permutations to align two transformers with different initializations that were trained on the same task. Mirzadeh et al. (2021) leveraged LMC for multitask and continual learning. Yunis et al. (2022) expanded LMC to more than two models, discovering a high-dimensional convex hull of low loss. Zhou et al. (2023) introduced Layerwise Linear Feature Connectivity (LLFC), demanding that the features of the models be linearly connected. Singh & Jaggi (2020); Akash et al. (2022) addressed a similar challenge of neuron alignment using optimal transport (Knight, 2008), and Liu et al. (2022) generalized it as a graph-matching task. Navon et al. (2023) solved the neuron alignments problem by training an equivariant deep weight space network. It is important to note that previous works which merged models from different initializations using permutations primarily focused on merging pairs of models trained on the same task, with many of them concentrating on MLPs and CNNs. In contrast, our method is capable of merging larger groups of models, with different initializations and tasks, and is specifically designed to handle transformers.
@@
Model merging technique (Goddard et al., 2024) has gained increasing interest in the past years, allowing the creation of stronger single-task and multi-task models. Akhlaghi & Sukhov (2018) showed that averaging the weights of multiple simple neural networks maintains their performance. Model soups (Wortsman et al., 2022) proposed averaging multiple models trained on the same task from identical initializations to enhance task accuracy. In addition, Weight averaging has been employed for various purposes, including improving optimization through checkpoint averaging (Izmailov et al., 2018), federated learning (McMahan et al., 2017), developing superior pre-trained models (Choshen et al., 2022; Don-Yehiya et al., 2022), improving Out-of-Distribution Generalization (Rame et al., 2022), achieving success in unseen tasks (Huang et al., 2023), and as augmentations for weight space networks (Shamsian et al., 2024). In many cases, spherical linear interpolation (SLERP) (Shoemake, 1985) is used to average the models’ weights. Matena & Raffel (2022) utilized Fisher-Weighted Averaging to fuse multiple models from the same initializations but trained on diverse tasks, resulting in a multi-task model. RegMean (Jin et al., 2023) addressed a similar scenario and proposed a closed-form solution that solves a local linear regression problem for each linear layer in the model. Ilharco et al. (2023) defined task vector by subtracting the parameters of a fine-tuned model from those of the pre-trained model, and used it to fuse models fine-tuned from the same pre-trained model. Yadav et al. (2024); Ortiz-Jimenez et al. (2024); Yang et al. (2024); Akiba et al. (2024) analyzed and proposed more merging methods based on task vectos. Sung et al. (2023) fused pairs of models trained on different modalities. ZipIt (Stoica et al., 2024) merged models from various initializations and tasks, focusing on MLPs and CNNs, by averaging pairs of highly correlated neurons between and within the models (see Appendix A.2 for more details). In contrast, our work focuses on MLPs and transformers, and can use much more complicated merging schemes.
@@
In Knowledge Distillation (KD), a small student model is trained to mimic the outputs of one or more larger teacher models (Ba & Caruana, 2014; Hinton et al., 2015; Stanton et al., 2021; Gou et al., 2021). Sau & Balasubramanian (2016) implemented a noise-based methodology to simulate learning a single task from multiple teachers. Tan et al. (2019); Khanuja et al. (2021) used several teachers, each translating between a specific language pair, to train a singular multilingual student. Clark et al. (2019); Park & Caragea (2023); Chelaramani et al. (2021) utilized multiple models and true labels to instruct a multi-task model. Pham et al. (2023) employed numerous quantized teachers trained on the same task to teach a quantized model, also leveraging true labels. Wu et al. (2021a) co-fine-tuned multiple teachers in downstream tasks with shared layers to instruct a student. Jacob et al. (2023) simultaneously trained multiple single-task models with a single multi-task student to facilitate the student’s optimization. Similar to our distillation baseline, Vongkulbhisal et al. (2019) utilized KD to unify knowledge from various models with different targets into a single model, relying solely on unlabeled data. Wu et al. (2021a); Zagoruyko & Komodakis (2017); Heo et al. (2019b; a) focused on using a single teacher’s outputs and inner features to train a single model. Park & Kwak (2019); Liu et al. (2020) applied multiple teachers trained on the same task, their inner features, and the true labels to train a single-task student. Inspired these works, we tried to merge transformers using a loss function that takes into account the inner features. However, this approach proved ineffective for both our method and distillation. See Appendix G.5 for details. Wan et al. (2024) generates next-token probability distributions from multiple LLMs on a shared corpus, merges them into a single fused probability matrix, and uses it to transfer knowledge into a single LLM. Crucially, their setting does not involve models trained on different tasks. In contrast, our setting involves classification models with a shared architecture but trained on different tasks. Therefore, using a shared corpus to generate a fused probability matrix would not enable effective knowledge merging in our case.
Adapter modules.
FS-Merge also resembles adapter-based methods (Houlsby et al., 2019; Chen et al., 2022; Zhang et al., 2024), which perform Parameter-Efficient Fine-Tuning (Hu et al., 2022; Zaken et al., 2021) of large models by injecting a small module ("adapters") into the original model. During fine-tuning, the original weights of the pre-trained model are frozen, and only the adapter weights are optimized. Similarly, FS-Merge injects new parameters into the model, which are the only ones optimized, though there are key differences. First, FS-Merge is used to merge the knowledge of multiple models, in contrast to adapter-based methods, which fine-tune a single model. Additionally, at the end of the training phase (or "merging"), the trainable parameters of FS-Merge are folded into the original weights, whereas in many adapter methods, the adapters remain in the fine-tuned models, adding complexity.
@@ A.2 Comparison to ZipIt
Our work strongly relates to ZipIt (Stoica et al., 2024), which also merges models from various initializations and tasks, with a focus on MLPs and CNNs. We will briefly explain the algorithm and its limitations. ZipIt uses the original training data to extract features from the models to be merged, $A$ and $B$. Then, for each layer $l$, it employs the features to identify pairs of highly correlated neurons using a greedy algorithm. Following this, it builds $M_{l}\in\mathbb{R}^{d_{l}\times 2\cdot d_{l}}$, which is zero except for entries corresponding to a matched pair $(i,j)$ indexed by $p$, where $M_{l}[p,i]=M_{l}[p,j]=\frac{1}{2}$. The purpose of this matrix is to merge features by averaging pairs of highly correlated neurons. Then, $U_{l}=2M_{l}^{\top}$ is used to reposition the merged features back to their original locations. After building $M_{l}$ and $U_{l}$, ZipIt proposes folding them with the original weights (Eq. 5) to create the weights of the merged model.
Like other current model merging methods (Ainsworth et al., 2023; Jin et al., 2023; Imfeld et al., 2023), ZipIt is efficient and focuses only on simple fusion schemes. It limit itself to pairing similar neurons, addressing only the permutation symmetries of neural networks (Hecht-Nielsen, 1990). Permutation symmetries mean that it is possible to swap any two neurons of a hidden layer in a neural network without altering its functionality. However, ZipIt falls short in handling the scale symmetries of neural networks (Neyshabur et al., 2015; Badrinarayanan et al., 2015; Phuong & Lampert, 2019), or in considering more complicated merging rules.
@@ B.1 Foldable SuperNet for Vision Transformers
Assuming there are two Vision Transformers (ViTs) (Dosovitskiy et al., 2021), trained on two distinct tasks, $A$ and $B$, we aim to define a Foldable SuperNet that combines the original weights of the ViTs with new learnable parameters $M,U$. This structure is designed so that all skip connections and layer normalizations (Ba et al., 2016; Xiong et al., 2020) operate on the merged dimension, and that there is a linear layer before or after every $M$ or $U$ matrix, allowing them to be folded after training. It is important to highlight that layer norms possess significantly fewer learnable parameters compared to other layers in ViTs. Therefore, we can initiate their parameters from a good starting point (e.g., the parameters of the first model to be merged) and proceed to optimize the parameters as usual, similar to strategies employed in previous merging works (Jordan et al., 2023; Stoica et al., 2024).
@@ B.1.1 Pre-processing
The tokens from both models are then concatenated to form $(Z^{A}||Z^{B})\in\mathbb{R}^{T\times 2\cdot d}$, and the Foldable SuperNet merges them using a learned matrix ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{\text{in}}}\in\mathbb{R}^{2\cdot d\times d}$ to produce $Z^{*}\in\mathbb{R}^{T\times d}$.
$Z^{*}=(Z^{A}||Z^{B}){\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{\text{in}}}\,.$
Subsequently, and like the ViT, the Foldable SuperNet applies layer normalization, with parameters ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma^{*}_{-1},\beta^{*}_{-1}}\in\mathbb{R}^{d}$, to generate the input for the transformer. These parameters will be optimized.
$\bar{X}_{0}^{*}=\text{LN}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma^{*}_{-1},\beta^{*}_{-1}}}(Z^{*})$
@@ B.1.2 The attention block
The Foldable SuperNet of the attention block at layer $l$ (Figure  3) receives merged features $X^{*}_{l}\in\mathbb{R}^{T\times d}$ as inputs, and applies layer normalization. It’s parameters ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma^{*}_{l},\beta^{*}_{l}}\in\mathbb{R}^{d}$ will be learned. Additionally, a learnable matrix ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}U_{l}}\in\mathbb{R}^{d\times 2\cdot d}$ is introduced to reconstruct the original features from the merged ones.
$\tilde{X}_{l}^{A}||\tilde{X}_{l}^{B}=\text{LN}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma^{*}_{l},\beta^{*}_{l}}}(X_{l}^{*}){\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}U_{l}}\,.$
@@
$Q_{l}^{*}=(\tilde{Q}^{A}_{l}||\tilde{Q}^{B}_{l}){\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{l}^{Q}}=(Q_{1,l}^{*},...,Q_{H,l}^{*})\,.$
The weights $W_{l}^{Q,A},W_{l}^{Q,B}\in\mathbb{R}^{d\times d}$ generate queries from the embeddings. The matrix ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{l}^{Q}}\in\mathbb{R}^{2\cdot d\times d}$ merges these features, which are then divided into $H$ heads, where each head $Q_{1,l}^{*}$ has dimensions $\mathbb{R}^{T\times\frac{d}{H}}$. Similarly, the keys and values are created by $W_{l}^{K},W_{l}^{V}\in\mathbb{R}^{d\times d}$, and compressed using the matrices ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{l}^{K},M_{l}^{V}}\in\mathbb{R}^{2\cdot d\times d}$ respectively.
Following this, and similar to the ViT, the Foldable SuperNet executes multi-head attention with the merged queries, keys and values and concatenates the features from the heads. In our Foldable SuperNet, this step also includes adding ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}U_{l}^{O}}\in\mathbb{R}^{d\times 2\cdot d}$ to reconstruct the original multi-head attention outputs. Then each ViT utilizes $W_{l}^{O}\in\mathbb{R}^{d\times d}$ to aggregate those outputs. We also use ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{l}^{O}}\in\mathbb{R}^{2\cdot d\times d}$ to compress it once more. This is followed by a skip connection.
$Y_{l}^{*}=X_{l}^{*}+\text{Concat}_{i}[...,\text{softmax}\left(\frac{Q_{i,l}^{*}{K_{i,l}^{*}}^{\top}}{\sqrt{d}}\right)V_{i,l}^{*},...]\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}U_{l}^{O}}\begin{pmatrix}W_{l}^{O,A}&0\\ 0&W_{l}^{O,B}\end{pmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{l}^{O}}\,.$